# Re-anchoring Quantum Monte Carlo with Tensor-Train Sketching

### v2

@@ 4. Convergence Analysis of cp-AFQMC
In this section, we provide a convergence analysis for cp-AFQMC to show how having a good trial wavefunction can improve the energy convergence. In Section 4.1, we first understand the effect of importance sampling in controlling the variance of the overlap of the walkers with the trial wavefunction in one step of cp-AFQMC. In Section 4.2, we then use this to show that a better trial wavefunction can lead to smaller energy variance. We further show that in cp-AFQMC, we can only hope for the convergence of energy estimator, but not the convergence of wavefunction.
@@ 4.1. Variance and bias per step of cp-AFQMC
In this section, we provide an explanation on how importance sampling controls the variance of the overlaps in one step of cp-AFQMC, and discuss the systematic bias caused by constrained-path approximation. We first consider a different representation of the walkers for the convenience of the analysis, where we define the following to be the walker that we keep track of:
@@
It might be the case where $\widetilde{P}_{k}^{(n)}(\vec{x})=0$ for some auxiliary fields $\vec{x}$. In this situation, some of the $\vec{x}$ values cannot be reached. Hence a systematic bias on $\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}$ is introduced, as explained in the following.
@@
In this case, the actual inner product $\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}$ is larger than $\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}$, thus a systematic bias towards the trial function $\Psi_{\mathrm{tr}}$ is introduced, which causes the sign problem.
@@ 4.2. Convergence analysis
In this section, we present how importance sampling with a trial wavefunction facilitates the convergence of energy. We focus on the case where there is no sign problem as in Lemma 2. We want to show that in such a benign situation, having a trial wavefunction that is closer to the ground truth wavefunction gives a smaller variance on the energy estimator (Section 4.2.1). We further show in Section 4.2.2 that unlike typical power method, we cannot hope for the walkers to converge to the ground-state wavefunction as the variance of the walkers is too large.
The results give us a better understanding of cp-AFQMC. Although it originates from the imaginary-time evolution (1), the wavefunction itself, represented as (5), does not necessarily converge to the ground-state wavefunction. However, we can still extract useful information out of it, for example, the ground-state energy.
@@
of which the second term $c_{\perp}\Psi_{\perp}$ represents the component of excited states, with $\Psi_{\perp}$ being a unit vector orthogonal to $\Psi_{0}$. The quantity $\left|c_{\perp}/c_{0}\right|$ then characterizes how close $\Psi_{\mathrm{tr}}$ is to the ground state $\Psi_{0}$. In what follows, we study how the convergence of the energy and the walkers depending on $\left|c_{\perp}/c_{0}\right|$.
@@ 4.2.1. Convergence analysis of the energy
In this section, we analyze how the mixed energy estimator converges to the ground-state energy in one step of cp-AFQMC. We want to show that the variance of the cp-AFQMC is controlled by $\left|c_{\perp}/c_{0}\right|$. When $\left|c_{\perp}/c_{0}\right|$ gets smaller, the variance of the energy gets smaller, i.e. the convergence of cp-AFQMC behaves like a vanilla power-iteration. In an ideal case where $\left|c_{\perp}/c_{0}\right|=0$, i.e. the trial function is exactly the ground-state wavefunction $\Psi_{\mathrm{tr}}=\Psi_{0}$, there is a zero-variance principle
We first state a lemma that characterizes the propagation of a walker when applying the exact operator $e^{-\Delta\tau H}$ instead of $\widetilde{B}_{k}^{(n)}(\vec{x})$:
@@
(24) $\left|\frac{\braket{\Psi_{\mathrm{tr}},He^{-\Delta\tau H}\Phi_{k}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}}-E_{0}\right|\leq e^{-\Delta\tau(E_{1}-E_{0})}\frac{1+a_{0}\tan\theta_{k}^{(n)}}{1-a_{0}\tan\theta_{k}^{(n)}}\frac{1+a_{1}\tan\beta_{k}^{(n)}}{1-a_{1}\tan\beta_{k}^{(n)}}\left|\frac{\braket{\Psi_{\mathrm{tr}},H\Phi_{k}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},\Phi_{k}^{(n)}}}-E_{0}\right|$
Proof.

### v4

@@ 4. Convergence Analysis of cp-AFQMC
In this section, we provide some convergence analysis for cp-AFQMC. In Section 4.1, we first understand the effect of importance sampling in controlling the variance of the overlap of the walkers with the trial wavefunction in one step of cp-AFQMC. In Section 4.2, we then use this to analyze the convergence of cp-AFQMC in terms of both the energy and the wavefunction itself. In Section 4.2.1, we prove that a better trial wavefunction can lead to a smaller variance in ground-state energy. This fact motivates our algorithm to obtain a high-quality trial wavefunction, so that we effectively reduce the variance in ground-state energy estimation within the cp-AFQMC procedure. In Section 4.2.2, we demonstrate that the random walkers individually exhibit a large variance, which cannot be eliminated by improving the trial wavefunction. Consequently, the finite ensemble of walkers (5) does not reliably converge to the ground-state wavefunction in practice. This observation further motivates our proposed algorithm: we extract an explicit wavefunction in TT form, approximating the ground-state wavefunction, from the noisy data produced by the cp-AFQMC procedure.
@@ 4.1. Variance and bias per step of cp-AFQMC
In this section, we provide an explanation on how importance sampling controls the variance of the overlaps in one step of cp-AFQMC, and discuss the systematic error caused by constrained-path approximation. We first consider a different representation of the walkers for the convenience of the analysis, where we define the following to be the walker that we keep track of:
@@
It might be the case where $\widetilde{P}_{k}^{(n)}(\vec{x})=0$ for some auxiliary fields $\vec{x}$. In this situation, some of the $\vec{x}$ values cannot be reached. Hence, a systematic error on $\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}$ is introduced, as explained in the following.
@@
In this case, the actual inner product $\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}$ is larger than $\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}$, thus a bias towards the trial function $\Psi_{\mathrm{tr}}$ is introduced, which causes the sign problem.
@@ 4.2. Convergence analysis
In this section, we present how importance sampling with a trial wavefunction facilitates the convergence of energy. We focus on the case where there is no sign problem as in Lemma 2. We want to show that in such a situation, having a trial wavefunction that is closer to the ground-state wavefunction gives a smaller variance on the energy estimator (Section 4.2.1). We further show in Section 4.2.2 that unlike typical power method, we cannot expect that the walkers converge to the ground-state wavefunction as the variance of the individual walkers, controlled primarily by the chosen form of $B(\vec{x})$, is too large.
The results give us a better understanding of cp-AFQMC. Although it originates from the imaginary-time evolution (1), the wavefunction itself, represented as (5), remains a statistical sampling of the ground-state wavefunction. In general, it does not provide a direct wavefunction of well-controlled variance [44]. However, we can still extract useful information out of it, for example, the ground-state energy.
@@
of which the second term $c_{\perp}\Psi_{\perp}$ represents the component of excited states, with $\Psi_{\perp}$ being a unit vector orthogonal to $\Psi_{0}$. The quantity $\left|c_{\perp}/c_{0}\right|$ then characterizes how close $\Psi_{\mathrm{tr}}$ is to the ground state $\Psi_{0}$. In what follows, we study how the convergence of the energy and the walkers depends on $\left|c_{\perp}/c_{0}\right|$.
@@ 4.2.1. Convergence analysis of the energy
In this section, we analyze how the mixed energy estimator converges to the ground-state energy in one step of cp-AFQMC. We want to show that the variance of cp-AFQMC is controlled by $\left|c_{\perp}/c_{0}\right|$. When $\left|c_{\perp}/c_{0}\right|$ gets smaller, the variance of the energy gets smaller, i.e. the convergence of cp-AFQMC behaves like a vanilla power iteration. In an ideal case where $\left|c_{\perp}/c_{0}\right|=0$, i.e. the trial function is exactly the ground-state wavefunction $\Psi_{\mathrm{tr}}=\Psi_{0}$, there is a zero-variance principle. Thus, we are motivated to improve the trial wavefunction in Algorithm 3 to reduce the variance in ground-state energy estimation.
We first state a lemma that characterizes the propagation of a walker when applying the exact operator $e^{-\Delta\tau H}$ instead of $\widetilde{B}_{k}^{(n)}(\vec{x})$.
@@
(25) $\left|\frac{\braket{\Psi_{\mathrm{tr}},He^{-\Delta\tau H}\Phi_{k}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}}-E_{0}\right|\leq e^{-\Delta\tau(E_{1}-E_{0})}\frac{1+a_{0}\tan\theta_{k}^{(n)}}{1-a_{0}\tan\theta_{k}^{(n)}}\frac{1+a_{1}\tan\beta_{k}^{(n)}}{1-a_{1}\tan\beta_{k}^{(n)}}\left|\frac{\braket{\Psi_{\mathrm{tr}},H\Phi_{k}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},\Phi_{k}^{(n)}}}-E_{0}\right|.$
With this result, we can then analyze how the energy changes when we apply the actual propagator $\widetilde{B}_{k}^{(n)}(\vec{x})$ to a walker $\Phi_{k}^{(n)}$.
Theorem 5.
Proof of both Lemma 4 and Theorem 5 can be found in Appendix B. We notice that the energy convergence consists of a deterministic power iteration part and a fluctuating part that is roughly $a_{0}\|H-E_{0}I\|_{2}\tan\theta_{k}^{(n)}=\left|c_{\perp}/c_{0}\right|\|H-E_{0}I\|_{2}\tan\theta_{k}^{(n)}$ (when $a_{0}$ is small). Here $\|H-E_{0}I\|_{2}$ resembles the energy scale, and $\tan\theta_{k}^{(n)}$ is a random variable that captures the deviation of a random walker from the ground state. While currently, we do not have a characterization of the variance of $\tan\theta_{k}^{(n)}$, if we assume that it has a bounded variance of $O(1)$, when $a_{0}$ gets small, i.e. the trial function gets closer to the ground-state wavefunction, the fluctuation term can be made small. In an ideal case where $a_{0}=0$ so that the trial function is exactly the ground-state wavefunction $\Psi_{\mathrm{tr}}=\Psi_{0}$, there is no fluctuation in the energy estimation. As a matter of fact, we have
$\frac{\braket{\Psi_{0},H\Psi}}{\braket{\Psi_{0},\Psi}}=E_{0},$
for any wavefunction $\Psi$.
4.2.2. Convergence analysis of the walkers
Now, we analyze the convergence of a particular walker $\Phi_{k}^{(n)}$ to the ground state, by keeping track of the angle between $\Phi_{k}^{(n)}$ and the ground state $\Psi_{0}$. As shown later in Theorem 6, unlike the results in energy convergence, even if the trial function is exactly the ground-state wavefunction, there is still an error term that does not vanish. In other words, the wavefunction represented by the ensemble of the walkers (5) has a large variance, and does not converge to the ground state reliably, regardless of the trial wavefunction. The convergence to the ground state is only in the sense of Monte Carlo sampling, as a statistical ensemble. Our algorithm addresses this limitation by reconstructing a wavefunction in TT form using information extracted from the random walkers.
Theorem 6.
Then,
$\tan\theta_{k}^{(n+1)}(\vec{x})\leq e^{-\Delta\tau(E_{1}-E_{0})}\tan\theta_{k}^{(n)}+e^{\Delta\tau E_{0}}\left(a_{0}\tan\theta_{k}^{(n)}+1\right)O\left(\frac{\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|}{|\braket{\Psi_{0},\Phi_{k}^{(n)}}|}\right)\\ \quad+a_{0}e^{2\Delta\tau E_{0}}O\left(\frac{\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|}{|\braket{\Psi_{0},\Phi_{k}^{(n)}}|}\right)^{2}.$
Proof of Theorem 6 can be found in Appendix B. As we can see, even when $a_{0}=0$, i.e. the trial wavefunction is exactly the ground-state wavefunction, there is still an error term $O\left(\frac{\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|}{|\braket{\Psi_{0},\Phi_{k}^{(n)}}|}\right)$. Thus, the individual walker is not guaranteed to get closer to the ground state. In fact, in general, one needs an exponential number of walkers to get $\frac{\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|}{|\braket{\Psi_{0},\Phi_{k}^{(n)}}|}<1$. This is due to the fact that, generally, one needs to sample an exponential number of $\widetilde{B}_{k}^{(n)}(\vec{x})$ to approximate $\exp(-\Delta\tau H)$. In other words, when sampling a multi-dimensional function, we can evaluate various properties with a polynomial cost, but to map out the function to pre-specified details by relying on the sample positions can require exponential cost.
5. Numerical Results
In this section, we apply our proposed algorithm on transverse-field Ising (TFI) systems [45, 46]. The Hamiltonian of the TFI model is given by
For more details about TFI system, we refer the reader to Appendix C. For the TFI system, the cp-AFQMC can be made free of sign problem, as analyzed in detail in Appendix C. Without a sign problem, the constrained-path approximation will not incur any systematic error, and the trial wavefunction $\Psi_{\mathrm{tr}}$ purely introduces importance sampling. Depending on the quality of $\Psi_{\mathrm{tr}}$, the efficiency as well as other controllable biases in the computation are affected, including those from fixed time step $\Delta\tau$ and finite population size $N$, which are discussed further in Appendix D. These provide an excellent test ground to illustrate and study our algorithm. As we show below, the algorithm leads to significant improvement across the board by allowing better importance sampling. Furthermore, by combining the features of AFQMC and DMRG, the algorithm shows advantages and extends the reach beyond each of its components separately.
5.1. cp-AFQMC-re-anchoring for transverse-field Ising model
In this section, we mainly compare our proposed algorithm, cp-AFQMC-re-anchoring, with vanilla cp-AFQMC for both 1D and 2D TFI systems, and illustrate that cp-AFQMC-re-anchoring significantly improves the accuracy and efficiency.
In all of our experiments, the step size of the imaginary-time evolution is set to $\Delta\tau=0.01$. Unless otherwise noted, we have verified that the remaining Trotter error is smaller than our statistical error bar or the signal for comparison. We set the initial trial wavefunction to be the fully disordered state
(28) $\Psi_{\mathrm{tr,0}}=\frac{1}{\sqrt{2^{d}}}\cdot\left(1,1\right)^{T}\otimes\cdots\otimes\left(1,1\right)^{T},$
which can also be interpreted as the eigenvector corresponding to the smallest eigenvalue of $H_{1}$ (The first term of the Hamiltonian in (27)). We implement TT-sketching to update the trial wavefunction every $50$ steps. At the $2000$-th step, we stop updating the trial wavefunction and keep using the latest one. We set the number of walkers to scale linearly with the number of spins. For systems with $16,32,64$ and $96$ spins, the numbers of walkers are set to $2000,4000,8000$ and $12000$, respectively.
In the TT-sketching procedure, we take $\delta=0.1$ in (16) for the sketching functions. For systems with $16$ or $32$ spins, the rank of the sketching functions is set to $60$. For systems with $64$ or $96$ spins, we set the rank to be $150$. The target rank of the sketched TT is set to $r=4$ in all experiments.
The 1D transverse-field Ising model with periodic boundary undergoes a quantum phase transition [48] at $g=1$ (see (27)), which is demonstrated in Figure 3 using the 1D system with $16$ spins as an example. To be more specific, We run our algorithm on the 1D system with $16$ spins with different transverse fields, and measure the ground-state energy $E_{0}(g)$. We also use Hellmann-Feynman theorem [49] to calculate the energy of the transverse field by
where we pick $\Delta g=0.01$ in practice.
(a) Ground-state energy
(b) Transverse-field energy
Figure 3. Results of cp-AFQMC-re-anchoring for $16$ spins in 1D with different transverse-field strength $g$. (a): ground-state energy; (b): transverse-field energy, where the critical point $g=1$ is illustrated with the black dashed line.
Since the quantum phase transition at the critical point $g=1$ makes it the most difficult case to deal with, we also compare cp-AFQMC-re-anchoring and the vanilla cp-AFQMC on larger systems, including 1D system with $d=32,64$ and $96$ spins, at $g=1$. The ground-state energy of 1D periodic system can be calculated analytically [50], and the ground-state wavefunction can be well approximated by density matrix renormalization group (DMRG) [51, 52].
In addition, we also test a quasi-1D system, where the spins are arranged on a $4\times 16$ lattice. We incorporate periodic boundary on the $4$-site axis, and open boundary on the $16$-site axis. Thus, it appears like a cylinder. The transverse field is also set to $g=1$. Despite the fact that the spins are arranged on a 2D lattice, the ground-state energy and ground-state wavefunction can still be well approximated by DMRG, due to the specific type of boundary conditions. Thus, we still use DMRG result as a reference.
As illustrated in Figure 4, for each system, the energy obtained by vanilla cp-AFQMC is of the order $O(10^{-3})$, while cp-AFQMC-re-anchoring can achieve the accuracy of order $O(10^{-5})$ or even smaller. The relative error of the energy is shown in Table 1. In addition, the overlap between our sketched trial wavefunction and the ground-state wavefunction is shown in Figure 5. For 1D systems, the sketched trial wavefunction in the form of a rank-$4$ TT can approximate the ground-state wavefunction quite well. For systems with $32$, $64$ and $96$ spins, the overlap can achieve around $0.9$, $0.8$, and $0.7$, respectively. For the $4\times 16$ cylinder system, the overlap converges to around $0.88$. As the quality of the trial wavefunction is improved, all sources of error in AFQMC will be reduced, resulting in smaller statistical errors in the estimators and reduced bias (population control and Trotter). This is clearly seen in the results in Table 1.
Figure 4. Comparison between the energy convergence of cp-AFQMC and cp-AFQMC-re-anchoring at $g=1$. From (a) to (d): $32$ spins in 1D; $64$ spins in 1D; $96$ spins in 1D; $4\times 16$ spins in 2D (cylinder).
Figure 5. The overlap between the ground-state wavefunction and the sketched wavefunction from cp-AFQMC-re-anchoring, in the form of a rank-$4$ TT. From (a) to (d): $32$ spins in 1D; $64$ spins in 1D; $96$ spins in 1D; $4\times 16$ spins in 2D (cylinder).
cp-AFQMC cp-AFQMC-re-anchoring
$32$ spins 1D $(+1.35\pm 0.31)\times 10^{-3}$ $(+0.44\pm 2.43)\times 10^{-5}$
$64$ spins 1D $(+1.88\pm 0.37)\times 10^{-3}$ $(+0.77\pm 1.07)\times 10^{-5}$
$96$ spins 1D $(+2.49\pm 0.33)\times 10^{-3}$ $(-4.95\pm 8.94)\times 10^{-6}$
$4\times 16$ spins 2D (cylinder) $(+4.18\pm 1.42)\times 10^{-3}$ $(+0.50\pm 1.46)\times 10^{-6}$
Table 1. Relative error of the energy. The new algorithm leads to a reduction of statistical error by more than two orders of magnitude, or $\mathcal{O}(10,000)$ reduction in computer time. Systematic errors are also reduced. The systematic errors still present in cp-AFQMC comes from the fixed imaginary-time step and the finite population size, which can be eliminated by extrapolating. See Appendix D for detailed discussion.
As an example of a 2D system, we test our algorithm on the $11\times 1$ 11 lattice with open boundary at the quantum critical point $g\approx 3.044$ [53]. In our experiments, we take $g=3$. The results are shown in Figure 6. For such a 2D system, there is no analytic solution for the ground-state energy anymore. A reference for the ground-state energy given by DMRG is $-3.17210\pm 1\times 10^{-5}$ [54], while the energy estimated by our method is $-3.17212\pm 4\times 10^{-5}$.
(a) Vanilla cp-AFQMC
(b) cp-AFQMC-re-anchoring
Figure 6. Energy convergence of (a) vanilla cp-AFQMC and (b) cp-AFQMC-re-anchoring for $11\times 1$ 11 system with open boundary at $g=3$. The reference ground-state energy, shown as the black dashed line, is obtained from DMRG [54].
5.2. Comparison with DMRG
Finally, we demonstrate that our algorithm allows the calculation of the ground-state energy to go beyond DMRG. The first comparison is conducted on the TFI system, where the spins are arranged on a $4\times 4$ lattice in 2D with a periodic boundary condition. Exact diagonalization can be performed for such a system, providing a reliable reference for comparison. We begin by focusing on the wavefunctions obtained from DMRG and cp-AFQMC-re-anchoring at various transverse-field strengths $g$. The TT-rank of the wavefunctions is set to $4$. As illustrated in Figure 7(a), the overlap between the DMRG wavefunction and the ground state exhibits a sharp transition as the transverse field $g$ changes. The transition occurs in the interval from $2.0$ to $2.6$. For $g$ larger than the transition interval, the DMRG wavefunction closely approximates the ground state, achieving an overlap of around $0.98$. However, when $g$ is smaller than the transition interval, DMRG does not recover the ground-state wavefunction, with an overlap only around $0.7$. Our algorithm effectively resolves this issue. The sketched wavefunction from the cp-AFQMC-re-anchoring procedure consistently recovers the ground state for various $g$, maintaining an overlap above $0.96$ throughout.
(a) Wavefunction overlap
(b) Energy estimation error
Figure 7. Comparison between cp-AFQMC-re-anchoring and DMRG for the $4\times 4$ system with periodic boundary with different transverse-field strengths. (a): comparison of the overlap between the ground-state wavefunction and the wavefunctions obtained by DMRG and TT-sketching; (b): comparison of the relative error in ground-state energy estimation by cp-AFQMC-DMRG and cp-AFQMC-re-anchoring.
Consistent with the observation above, the computed ground-state energies also show an improvement with the algorithm, as illustrated in Figure 7(b). We take a closer look at $g=2.0$, where the two algorithms exhibit a significant difference. The overlap between the ground-state and the sketched wavefunction during the cp-AFQMC-re-anchoring procedure is shown in Figure 8(a). In terms of the ground-state energy, Figure 8(b) clearly shows that cp-AFQMC-re-anchoring provides a more accurate estimation compared to cp-AFQMC-DMRG. Specifically, the relative energy error for cp-AFQMC-re-anchoring is $(0.59\pm 1.57)\times 10^{-5}$, while the error for cp-AFQMC-DMRG is $(+1.87\pm 0.08)\times 10^{-4}$, which exhibits a substantial bias.
(a) Wavefunction overlap
(b) Energy convergence
Figure 8. Comparison between cp-AFQMC-re-anchoring and cp-AFQMC-DMRG for the $4\times 4$ system with periodic boundary at $g=2.0$. (a): comparison of the overlap between the ground-state wavefunction and the wavefunctions obtained by both methods, in the form of a rank-$4$ TT; (b): comparison of the energy convergence.
Apart from ground-state energy, we can also assess the advantages of our algorithm in the estimation of other observables. To estimate the expectation of an arbitrary observable $O$ in the ground state
$\braket{O}_{\mathrm{gs}}=\frac{\braket{\Psi_{0},O\Psi_{0}}}{\braket{\Psi_{0},\Psi_{0}}},$
we apply the extrapolated estimator [55]
$\braket{O}_{\mathrm{extrap}}=2\braket{O}_{\mathrm{mix}}-\braket{O}_{\mathrm{tr}},$
where $\braket{O}_{\mathrm{mix}}$ denotes the mixed estimator, analogous to (13), and
$\braket{O}_{\mathrm{tr}}=\frac{\braket{\Psi_{\mathrm{tr}},O\Psi_{\mathrm{tr}}}}{\braket{\Psi_{\mathrm{tr}},\Psi_{\mathrm{tr}}}}.$
(Because the quality of our trial wavefunction is very high, the extrapolated estimator is sufficient to remove any residual systematic error.) As an example, we study the average spin in $z$-direction:
(29) $\mathcal{S}_{z}=\frac{1}{d}\sum_{i=1}^{d}\sigma_{i}^{z}.$
Due to the symmetry of the Hamiltonian (27) under spin-flip transformations in $\sigma_{z}$, specifically the invariance of the Hamiltonian when all $\sigma_{i}^{z}$ are flipped to $-\sigma_{i}^{z}$, the expectation value of $\mathcal{S}_{z}$ in the ground state is $\braket{\mathcal{S}_{z}}_{\mathrm{gs}}=0$. Numerical results in Figure 9 demonstrate that cp-AFQMC-re-anchoring preserves the $\sigma_{z}$ symmetry, which is broken in cp-AFQMC-DMRG.
Figure 9. Average spin in $z$-direction (defined in (29)) of the $4\times 4$ system with periodic boundary at $g=2.0$.
Another example is the spin-correlation. For two spins $i$ and $j$, their correlation is calculated as
(30) $\mathrm{Cor}(i,j)=\frac{\mathrm{Cov}(\sigma_{i}^{z},\sigma_{j}^{z})}{\sqrt{\mathrm{Var}(\sigma_{i}^{z})\mathrm{Var}(\sigma_{j}^{z})}}=\frac{\braket{\sigma_{i}^{z}\sigma_{j}^{z}}-\braket{\sigma_{i}^{z}}\braket{\sigma_{j}^{z}}}{\sqrt{(1-\braket{\sigma_{i}^{z}}^{2})(1-\braket{\sigma_{j}^{z}}^{2}})},$
where we have incorporated the fact that $(\sigma_{i}^{z})^{2}=(\sigma_{j}^{z})^{2}=I$.
Under periodic boundary conditions, the correlation between two spins depends solely on their Taxicab distance [56], which is simply the number of steps needed to reach one spin from another, as illustrated in Figure 10(a). We select spin $1$ as the reference (highlighted in red) and monitor its correlation with the spins labeled in green along the arrows. The comparison of our algorithm and cp-AFQMC-DMRG is presented in Figure 10(b). It is seen that cp-AFQMC-re-anchoring recovers the correct spin-correlation perfectly, whereas cp-AFQMC-DMRG does not.
(a) Labels of the spins
(b) Spin-correlation
Figure 10. Spin-correlation (defined in (30)) of the $4\times 4$ system with periodic boundary at $g=2.0$. (a): labeling of the spins; (b): correlations with the reference spin $1$.
As a further illustration, we test both algorithms on a larger system with 64 spins arranged on an $8\times 8$ lattice with periodic boundary, and estimate the ground-state energy of this system. The transverse field is set to $g=3.0$. For such a large system, exact diagonalization is no longer feasible, so the reference for the ground-state energy is obtained using DMRG with increasing ranks. As shown in Figure 11, the ground-state energy estimated by cp-AFQMC-re-anchoring matches the DMRG results obtained with a very high rank TT. The energy convergence plots for both cp-AFQMC-re-anchoring and cp-AFQMC-DMRG are presented in Figure 11. In both cases, the rank of the trial wavefunctions is set to $4$. As can be seen, the energy estimated by cp-AFQMC-DMRG exhibits a larger error compared to that of our algorithm.
Figure 11. Results for the $8\times 8$ system with periodic boundary at $g=3.0$. The rank of the trial wavefunctions in cp-AFQMC is set to $4$. (a): comparison between the ground-state energy estimated by DMRG with increasing ranks and cp-AFQMC-re-anchoring; (b): comparison between the energy convergence of cp-AFQMC-re-anchoring and cp-AFQMC-DMRG.
6. Conclusion
In this paper, we proposed a framework for studying the ground-state properties of quantum many-body systems. Our algorithm improves cp-AFQMC by periodically updating the trial wavefunction using TT-sketching. Systematic analysis of the algorithm and detailed numerical tests on the transverse-field Ising model are presented. The algorithm combines key features of AFQMC with TT/MPS, providing a route for a more powerful approach that extends the capabilities of each. This work opens a number of future directions, including extension and application of this framework to electronic systems.
Acknowledgments
YK is partially funded by DMS-2339439, DE-SC0022232, and Sloan Research Fellowship.
The authors would like to thank Miles Stoudenmire for helpful discussions on the use of the ITensor library in our numerical simulations.
Appendix A Time Complexity of TT Calculation
In this section, we investigate numerically the time complexity of tensor-train calculations. Specifically, we verify that computing the inner product between a TT trial wavefunction and the walkers scales linearly with both the number of spins and the number of walkers, as shown in Figure 12. Similarly, the computational cost of TT-sketching has a linear scaling in both parameters, as illustrated in Figure 13.
(a) Scaling with number of spins
(b) Scaling with number of walkers
Figure 12. Computational cost of calculating the inner product of a TT trial wavefunction and the walkers. (a): different number of spins with $2000$ walkers; (b): different number of walkers for $16$ spins.
(a) Scaling with number of spins
(b) Scaling with number of walkers
Figure 13. Computational cost of TT-sketching. (a): different number of spins with $2000$ walkers; (b): different number of walkers for $16$ spins.
Appendix B Details on Convergence Analysis of cp-AFQMC
In this section, we provide proofs for Lemma 4, Theorem 5 and Theorem 6.
Proof of Lemma 4.